# Faster Reinforcement Learning by Freezing Slow States

### v2

@@ Abstract
We study infinite horizon Markov decision processes (MDPs) with fast-slow structure, where some state variables evolve rapidly (“fast states”) while others change more gradually (“slow states”). Such structure is common in real-world problems where sequential decisions need to be made at high frequencies over long horizons, where slowly evolving information also influences optimal decisions. Examples include inventory control under slowly changing demand, or dynamic pricing with gradually shifting consumer behavior. Modeling the problem at the natural decision frequency leads to MDPs with discount factors close to one, making them computationally challenging. We propose a novel approximation strategy that “freezes” slow states during a phase of lower-level planning, solving finite-horizon MDPs conditioned on a fixed slow state, and then applying value iteration to an auxiliary upper-level MDP that evolves on a slower timescale. Freezing states for short periods of time leads to easier-to-solve lower-level problems, while a slower upper-level timescale allows for a more favorable discount factor. On the theoretical side, we analyze the regret incurred by our frozen-state approach, which leads to simple insights on how to trade off computational budget versus regret. Empirically, we demonstrate that frozen-state methods produce high-quality policies with significantly less computation, and we show that simply omitting slow states is often a poor heuristic.
@@ Introduction
Markov decision processes (MDPs) offer a foundational framework for modeling sequential decision making under uncertainty. In many real-world applications, decisions must be made at high frequencies over extended time horizons. From a modeling standpoint, this translates into MDPs with discount factors close to one, capturing the long-term nature of the objective. However, such long-horizon problems can pose computational challenges: standard algorithms like value iteration (VI), which is a core building block in many reinforcement learning (RL) algorithms, rely on the contraction property of the Bellman operator to guarantee convergence. As the discount factor approaches one, this contraction weakens, leading to slow convergence and degraded performance within a given computational budget (Bertsekas and Tsitsiklis, 1996a; Jiang et al., 2015).
In a wide range of applications, state variables naturally evolve at different timescales. For example, in inventory control problems, the inventory level may fluctuate rapidly from one period to the next, while exogenous factors that influence demand may evolve more gradually. In dynamic pricing problems, the (endogenous) price history can change frequently, but user behavior shifts more slowly. We refer to this as fast-slow structure, where the state decomposes into fast-evolving and slow-evolving components. Although slow states may evolve gradually, they often encode critical contextual information and omitting them can result in significantly suboptimal policies and high regret.
Solving MDPs that fully capture both fast and slow dynamics at the natural decision frequency typically requires long effective horizons, posing serious computational burdens. To manage complexity, practitioners often resort to simplifying assumptions, such as fixing or omitting slow states altogether. While this reduces computational burden, it can degrade policy quality by discarding important structure. This paper introduces and formalizes the notion of fast-slow MDPs, and studies the computational implications of a novel multi-scale approximation strategy called frozen-state value iteration (FSVI).
@@
In this paper, we propose an approximation strategy that uses two levels of planning, aiming to provide a middle ground between solving the full MDP and completely ignoring slow states. We temporarily “freeze” the slow state during lower-level planning—solving finite-horizon MDPs conditioned on a fixed slow state—and then apply value iteration to an auxiliary upper-level MDP that evolves on a slower timescale. This approach simplifies the lower-level problems, making them more tractable, while the slower upper-level dynamics enable the use of a more favorable (i.e., smaller) effective discount factor.Specifically, we make the following contributions:
@@
    We propose a frozen-state approximation to the reformulated MDP, along with an associated frozen-state value iteration (FSVI) algorithm, where the slow state is frozen in the lower-level problem, while each period in the upper-level problem “releases” the slow state. Computational benefits arise in several ways: (1) frozen states simplify the dynamics of the lower-level MDP (fewer successor states), (2) more favorable discount factors in the upper-level problem through reuse of the lower-level policy, which is computed only once, (3) the lower-level MDP thus becomes separable into independent MDPs, opening the door to speedups via parallel computation. Solving the frozen-state approximation gives a policy that switches between one action from upper-level policy and $T-1$ actions from the lower-level policy. We give a theoretical analysis that upper bounds the expected regret from applying this policy compared to the optimal policy.
@@
    We adapt our frozen-state approach to a more practical reinforcement learning setting where a generative model (or simulator) is available but the MDP is not fully known. Specifically, we develop frozen-state fitted value iteration (FSFVI), a version of fitted value iteration (Munos and Szepesvári, 2008) that incorporates our two-level planning strategy. FSFVI can learn effective policies from sampled transitions, making our method applicable to model-free or simulation-based environments. We also provide a theoretical regret bound for a special case of FSFVI using linear function approximation (see Appendix B).
@@
    Lastly, we perform a systematic empirical study on three problem settings: (1) gridworld with spatial tasks, (2) inventory control with fixed costs, and (3) dynamic pricing with reference effects. We show that the proposed algorithms based on the frozen-state approximation quickly converge to good policies using significantly less computation compared to standard methods. Notably, our results show that ignoring the slow state leads particularly poor results.
@@
Note that since $J_{T}^{*}\equiv 0$, it follows that $J_{1}^{\tilde{\bm{\pi}}^{*}}=J_{1}^{*}$, meaning that we can use $J_{1}^{*}$ to represent the value of the lower-level problem.
@@
What about when $J_{T}^{*}\neq 0$? In some applications, it may be natural or desirable to assign a non-zero terminal value at time $T$. This would allow for a heuristic continuation value based on the algorithm designer’s domain knowledge to be incorporated into the frozen-state approximation. In this case, the lower-level policy $\tilde{\bm{\pi}}^{*}$ can still be computed using finite-horizon dynamic programming, starting from the non-zero $J_{T}^{*}$ as the terminal condition in the recursion (10). However, for consistency with the structure of the overall frozen-state approximation—particularly its interpretation as approximating a subproblem within a larger hierarchical MDP—it is important to evaluate the performance of the resulting policy $\tilde{\bm{\pi}}^{*}$ using the original objective (9). In other words, when $J_{T}^{*}\neq 0$, $J_{1}^{\tilde{\bm{\pi}}^{*}}$ is no longer guaranteed to be the same as $J_{1}^{*}$. Therefore, to use a non-zero terminal value, we need to perform an additional policy evaluation step for $\tilde{\bm{\pi}}^{*}$ to obtain $J_{1}^{\tilde{\bm{\pi}}^{*}}$. This is a relatively simple one-time calculation that can be performed concurrently with the computation of $\tilde{\bm{\pi}}^{*}$. We assume throughout the paper that $J_{T}^{*}\equiv 0$ for simplicity.
@@
Throughout the paper, we use the notation $V^{\mu}(J_{1},\bm{\pi})$ and $V^{*}(J_{1},\bm{\pi})$ to refer to the value function (i.e., $\mathcal{S}\rightarrow\mathbb{R}$) obtained when the MDP is evaluated or solved for a fixed $J_{1}$ and $\bm{\pi}$. We also define the Bellman operator associated with (14):
@@
In this section, we introduce the our new approach: the frozen-state value iteration (FSVI) algorithm. We start by mentioning that the naive approach to solving the base model MDP (5) is to directly apply standard VI (see, e.g., Bertsekas and Tsitsiklis (1996b)). For completeness, we provide the full description in Algorithm 1. The main ideas of our approach are:
@@
    Solve the lower-level MDP with frozen states to obtain a policy $\tilde{\bm{\pi}}^{*}$ and its value $J_{1}^{*}$. Since the lower-level problem is a finite horizon MDP, it can be solved exactly using $T-1$ steps of VI.
@@
    Apply value iteration (VI) to the upper-level problem starting with some initial value function $V^{0}$, while using $J_{1}^{*}$ to approximate the $T$-horizon reward and $\tilde{\bm{\pi}}^{*}$ for $T$-step transitions. Note that this is an infinite-horizon MDP that operates at a slower timescale and enjoys a much more favorable discount factor of $\gamma^{T}$.
A detailed specification is given in Algorithm 2. We denote the resulting value function approximation after $k$ iterations of value iteration as $V^{k}$, from which we obtain a policy $\tilde{\mu}^{k}$. The $T$-periodic policy output by FSVI is formed by combining $\tilde{\mu}^{k}$ with the optimal finite-horizon policy $\tilde{\bm{\pi}}^{*}$ from the lower-level MDP: $(\tilde{\mu}^{k},\tilde{\bm{\pi}}^{*})$.
@@
Algorithm 1 Exact VI for the Base Model
@@
    Problems with implicit hierarchical structure. Many real-world applications do not exhibit explicit hierarchy in their MDP formulations but nevertheless contain natural two-level behavioral structure. We illustrate this type of problem in Section 9 through our gridworld with spatial tasks domain: the agent first “accepts” a task (each task comes with a starting and ending location) to undertake and then executes a sequence of navigation actions to complete it. This problem was motivated by operations domains such as warehouse robotics (Yang et al., 2020) and on-demand spatial services from the driver’s perspective (Chung et al., 2018; Jiang et al., 2020; Ong et al., 2021; Ashkrof et al., 2024), where high-level task acceptance decisions are followed by fast-paced local execution.
@@
    Problems with natural cyclical behavior. In some domains, optimal or near-optimal behavior follows a cyclic structure. A classic example is inventory control, where the policy repeatedly builds up and depletes inventory. We show results on inventory control with fixed costs in Section 9. Related problems, such as energy or commodity trading (Kim and Powell, 2011; Carmona and Ludkovski, 2010; Löhndorf and Minner, 2010; Jiang and Powell, 2015) or inventory repositioning (He et al., 2020; Benjaafar et al., 2022), may also exhibit similar cyclical behavior. Another domain we consider in Section 9 is a dynamic pricing problem with reference effects, where cyclic pricing strategies have been shown to perform well under various assumptions (Wang, 2016; Chen et al., 2017). In all of the above settings, freezing the slow state over a short horizon can capture the cyclical dynamics of the optimal policy.
@@
Let us now discuss the computational cost of FSVI. Here, we take a “dynamic programming” perspective, [^6] In our numerical results of Section 9, we will take a different, more practical perspective, where we work with a generative model or simulator of the environment. where the model is known (and expectations can be computed). It is well-known that each iteration of standard VI has time complexity $\mathcal{O}\bigl{(}|\mathcal{S}|^{2}|\mathcal{A}|\bigr{)}$, which provides a contraction factor of $\gamma$ (Littman et al., 1995). The upper level of FSVI, on the other hand, enjoys an improved contraction factor $\gamma^{T}$ with the same per-iteration running time of $\mathcal{O}\bigl{(}|\mathcal{S}|^{2}|\mathcal{A}|\bigr{)}$, given that we pay a one-time fixed cost of solving the lower level. The $\mathcal{O}\bigl{(}|\mathcal{S}|^{2}|\mathcal{A}|\bigr{)}$ consists of $|\mathcal{S}||\mathcal{A}|$ due to the number of state-action pairs at which to compute the Bellman update and another factor of $|\mathcal{S}|$ due to the number of successor states. Since freezing slow states restricts the successor states to $\mathcal{Y}$, each iteration of the lower-level VI (Lines 2-7 of Algorithm 2) has running time $\mathcal{O}\bigl{(}|\mathcal{X}||\mathcal{Y}|^{2}|\mathcal{A}|\bigr{)}$. An additional $\mathcal{O}\bigl{(}|\mathcal{S}|^{2}\,T\bigr{)}$ is required to compute the $T$-step transition probabilities of following $\tilde{\bm{\pi}}^{*}$, to be used in the upper-level VI, resulting in a one-time fixed cost of $\mathcal{O}(|\mathcal{X}||\mathcal{Y}|^{2}|\mathcal{A}|\,T+|\mathcal{S}|^{2}\,T)$. Particularly when $|\mathcal{X}|$ is large, this can be a reasonable fixed cost to pay in order to get the much improved discount factor of $\gamma^{T}$ going forward.
@@
The detailed proof is in Appendix C.3. ∎
@@
Regret of the Ideal Frozen-State Policy $(\tilde{\mu}^{*},\tilde{\bm{\pi}}^{*})$
In this section, we will show an upper bound on the regret from applying the $T$-periodic policy $(\tilde{\mu}^{*},\tilde{\bm{\pi}}^{*})$ instead of the optimal policy $\nu^{*}$ in the base model. Note that by “ideal,” we mean that this is the policy obtained if we were able to perfectly solve the frozen-state approximation. We start with definitions of the regret of both stationary and $T$-periodic policies.
@@
In this section, we derive a bound on $\mathcal{R}(\tilde{\mu}^{*},\tilde{\bm{\pi}}^{*})$, the regret of applying $T$-periodic policy $(\tilde{\mu}^{*},\tilde{\bm{\pi}}^{*})$ to the base model. First, we start with a general lemma, that will be used throughout the paper as a tool to analyze variants of FSVI.

### v3

@@ Abstract
We study infinite horizon Markov decision processes (MDPs) with fast-slow structure, where some state variables evolve rapidly (“fast states”) while others change more gradually (“slow states”). This structure commonly arises in practice when decisions must be made at high frequencies over long horizons, and where slowly changing information still plays a critical role in determining optimal actions. Examples include inventory control under slowly changing demand indicators or dynamic pricing with gradually shifting consumer behavior. Modeling the problem at the natural decision frequency leads to MDPs with discount factors close to one, making them computationally challenging. We propose a novel approximation strategy that “freezes” slow states during phases of lower-level planning and subsequently applies value iteration to an auxiliary upper-level MDP that evolves on a slower timescale. Freezing states for short periods of time leads to easier-to-solve lower-level problems, while a slower upper-level timescale allows for a more favorable discount factor. On the theoretical side, we analyze the regret incurred by our frozen-state approach, which leads to simple insights on how to trade off regret versus computational cost. Empirically, we benchmark our new frozen-state methods on three domains, (i) inventory control with fixed order costs, (ii) a gridworld problem with spatial tasks, and (iii) dynamic pricing with reference-price effects. We demonstrate that the new methods produce high-quality policies with significantly less computation, and we show that simply omitting slow states is often a poor heuristic.
@@ Introduction
Markov decision processes (MDPs) offer a foundational framework for modeling sequential decision making under uncertainty. In many real-world applications, decisions must be made at high frequencies over extended time horizons. From a modeling standpoint, this translates into MDPs with discount factors close to one (which capturing the long-term nature of the objective). However, such long-horizon problems can pose computational challenges: standard algorithms like value iteration (VI), which is a core building block in many reinforcement learning (RL) algorithms, rely on the contraction property of the Bellman operator to guarantee convergence. As the discount factor approaches one, this contraction weakens, leading to slow convergence and degraded performance within a given computational budget (Bertsekas and Tsitsiklis, 1996; Jiang et al., 2015).
In a wide range of applications, state variables naturally evolve at different timescales. For example, in inventory control problems, the inventory level may fluctuate rapidly from one period to the next, while exogenous factors that influence demand may evolve more gradually. In dynamic pricing problems, the (endogenous) price history can change frequently, but user behavior shifts more slowly. We refer to this as fast-slow structure, where the state decomposes into fast-evolving and slow-evolving components. Although slow states may evolve gradually, they often encode critical contextual information, and omitting them can result in suboptimal policies and high regret.
Solving MDPs that fully capture both fast and slow dynamics at the natural decision frequency typically requires long effective horizons, which can pose serious computational burdens. To manage complexity, practitioners often resort to simplifying assumptions, such as fixing or omitting slow states altogether. While this reduces computational burden, it can degrade policy quality by discarding important information. This paper introduces and formalizes the notion of fast-slow MDPs, and studies the computational implications of a novel multi-scale approximation strategy called frozen-state value iteration (FSVI).
@@
In this paper, we propose an approximation strategy that uses two levels of planning, aiming to provide a middle ground between solving the full MDP and completely ignoring slow states. In our approach, we first temporarily “freeze” the slow state during lower-level planning. Concretely, we compute policies based on finite-horizon MDPs that are conditioned on a fixed slow state. Subsequently, we apply value iteration to an auxiliary upper-level MDP that evolves on a slower timescale by leveraging the lower-level policies. This approach makes the lower-level problems more tractable, while the slower upper-level dynamics enable the use of a more favorable (i.e., smaller) effective discount factor. Specifically, we make the following contributions:
@@
    We propose a frozen-state approximation to the reformulated MDP, along with an associated frozen-state value iteration (FSVI) algorithm, where the slow state is frozen in the lower-level problem, while each period in the upper-level problem “releases” the slow state. Computational benefits arise in several ways: (1) frozen states simplify the dynamics of the lower-level MDP (fewer successor states), (2) the discount factor in the upper-level problem is more favorable due to the reuse of the lower-level policy, which is computed only once, (3) the lower-level MDP thus becomes separable into independent MDPs, opening the door to speedups via parallel computation. Solving the frozen-state approximation gives a policy that switches between one action from upper-level policy and $T-1$ actions from the lower-level policy. We give a theoretical analysis that upper bounds the expected regret from applying this policy compared to the optimal policy.
@@
    We adapt our frozen-state approach to a more practical reinforcement learning setting where a generative model (or simulator) is available but the MDP is not fully known. Specifically, we develop frozen-state fitted value iteration (FSFVI), a version of fitted value iteration (Munos and Szepesvári, 2008) that incorporates our two-level planning strategy. FSFVI can learn effective policies from sampled transitions, making our method applicable to model-free or simulation-based environments. We also provide a theoretical regret bound for a special case of FSFVI using linear function approximation (see Appendix C).
@@
    Lastly, we perform a systematic empirical study on three problem settings: (1) inventory control with fixed costs, (2) gridworld with spatial tasks, and (3) dynamic pricing with reference effects. We show that our proposed algorithms (based on the frozen-state approximation) quickly converge to good policies using significantly less computation compared to standard methods. Notably, our results show that ignoring the slow state, i.e., modeling the system as if the true state variable only includes the fast component, [^1] See Appendix B.1 for details on how we simulate this behavior in our experiments. leads to poor results.
@@
Finally, our work contributes to the broader literature on leveraging MDP structure to improve reinforcement learning and approximate dynamic programming algorithm. Prior research has exploited structural properties of the value function, such as convexity (Pereira and Pinto, 1991; Philpott and Guan, 2008; Nascimento and Powell, 2009; Benjaafar et al., 2022) and monotonicity (Papadaki and Powell, 2002; Jiang and Powell, 2015a), as well as structural features of the optimal policy (Kunnumkal and Topaloglu, 2008). Additional examples include factored MDPs (Osband and Van Roy, 2014) and weakly-coupled systems (Killian et al., 2021; El Shar and Jiang, 2023; Brown and Smith, 2025; Nadarajah and Cire, 2025). Our contribution to this line of work lies in identifying a new form of MDP structure, which then informs the design of a new algorithmic approach.
@@
$H$ Bellman operator for the base MDP
$\tilde{H}$ Bellman operator for lower-level frozen-state problem (discount factor $\gamma$)
$F_{J_{1},\bm{\pi}}$ Bellman operator for upper-level problem given lower-level $J_{1}$ and $\bm{\pi}$ (discount factor $\gamma^{T}$)
@@
Note that using $J_{T}^{*}\equiv 0$ is aligned with the fact that the lower-level problem (9) contains $T-1$ periods. It thus follows that $J_{1}^{\tilde{\bm{\pi}}^{*}}=J_{1}^{*}$, meaning that we can use $J_{1}^{*}$ to represent the value of the lower-level problem. We will see in the next section that the upper-level problem depends on an approximation of $J_{1}^{\tilde{\bm{\pi}}^{*}}$, which is the value of the lower-level policy in $T-1$ periods (and zero terminal value after that).
@@
What about when one wants to solve the lower-level problem with some $J_{T}^{*}\neq 0$? In some applications, it may be natural or desirable to assign a non-zero terminal value at time $T$. This would allow for a heuristic continuation value based on the algorithm designer’s domain knowledge. In this case, the lower-level policy $\tilde{\bm{\pi}}^{*}$ can still be computed using finite-horizon dynamic programming, starting from the non-zero $J_{T}^{*}$ as the terminal condition in the recursion (10). However, for consistency of the upper-level problem, it is important to evaluate the performance of the resulting policy $\tilde{\bm{\pi}}^{*}$ using the original objective (9) (without the terminal value). In other words, when $J_{T}^{*}\neq 0$, $J_{1}^{\tilde{\bm{\pi}}^{*}}$ is no longer guaranteed to be the same as $J_{1}^{*}$. Therefore, to use a non-zero terminal value, we need to perform an additional policy evaluation step for $\tilde{\bm{\pi}}^{*}$ to obtain $J_{1}^{\tilde{\bm{\pi}}^{*}}$. This is a relatively simple one-time calculation that can be performed concurrently with the computation of $\tilde{\bm{\pi}}^{*}$. We assume throughout the paper that $J_{T}^{*}\equiv 0$ for simplicity.
@@
Throughout the paper, we use the notations $V^{\mu}(J_{1},\bm{\pi}):\mathcal{S}\rightarrow\mathbb{R}$ and $V^{*}(J_{1},\bm{\pi}):\mathcal{S}\rightarrow\mathbb{R}$ to refer to the value function obtained when the MDP is evaluated or solved for a fixed $J_{1}$ and $\bm{\pi}$, respectively. We also define the Bellman operator associated with (14):
@@
In this section, we introduce the our new approach: the frozen-state value iteration (FSVI) algorithm. We start by mentioning that the naive approach to solving the base model MDP (5) is to directly apply standard VI (see, e.g., Bertsekas and Tsitsiklis (1996)). For completeness, we provide the full description of standard VI in Algorithm 1.
Our new approach, FSVI, is given in Algorithm 2. The main ideas of the algorithm are:
@@
    Solve the lower-level MDP with frozen states to obtain a policy $\tilde{\bm{\pi}}^{*}$ and its value $J_{1}^{*}$. Since the lower-level problem is a finite horizon MDP, it can be solved exactly using $T-1$ steps of VI. An alternative way of viewing the lower-level problem is that it is $|\mathcal{X}|$ independent MDPs, which we can potentially solve in parallel. Note that Algorithm 2 emphasizes this by looping over the slow state $x\in\mathcal{X}$ in Line 1.
@@
    Apply value iteration (VI) to the upper-level problem starting with some initial value function $V^{0}$, while using $J_{1}^{*}$ to approximate the $T$-horizon reward and $\tilde{\bm{\pi}}^{*}$ for $T$-step transitions. Note that this is an infinite-horizon MDP that operates at a slower timescale and enjoys a more favorable discount factor of $\gamma^{T}$.
We denote the resulting value function approximation after $k$ iterations of value iteration as $V^{k}$, from which we obtain a policy $\tilde{\mu}^{k}$. The $T$-periodic policy output by FSVI is formed by combining $\tilde{\mu}^{k}$ with the optimal finite-horizon policy $\tilde{\bm{\pi}}^{*}$ from the lower-level MDP: $(\tilde{\mu}^{k},\tilde{\bm{\pi}}^{*})$.
@@
Algorithm 1 Exact VI for the Base Model (Base VI)
@@
    Problems with implicit hierarchical structure. Many real-world applications do not exhibit explicit hierarchy in their MDP formulations but nevertheless contain natural two-level behavioral structure. We illustrate this type of problem in Section 9 through our gridworld with spatial tasks domain: the agent first “accepts” a task (each task comes with a starting and ending location) to undertake and then executes a sequence of navigation actions to complete it. This problem was motivated by operations domains such as warehouse robotics (Yang et al., 2020) and on-demand spatial services from the driver’s perspective (Chung et al., 2018; Jiang et al., 2020; Ong et al., 2021; Ashkrof et al., 2024), where high-level task acceptance decisions are followed by fast-paced and relatively more myopic execution.
@@
    Problems with natural cyclical behavior. In some domains, optimal or near-optimal behavior follows a cyclic structure. A classic example is inventory control, where the policy repeatedly builds up and depletes inventory. We show results on inventory control with fixed costs in Section 9. Related problems, such as energy or commodity trading (Kim and Powell, 2011; Carmona and Ludkovski, 2010; Löhndorf and Minner, 2010; Jiang and Powell, 2015b) or inventory repositioning (He et al., 2020; Benjaafar et al., 2022), may also exhibit similar cyclical behavior. Another domain we consider in Section 9 is a dynamic pricing problem with reference effects, where cyclic pricing strategies have been shown to perform well under various assumptions (Wang, 2016; Chen et al., 2017). In all of the above settings, freezing the slow state over a short horizon can capture the cyclical dynamics of the optimal policy.
@@
Let us now discuss the computational cost of FSVI. Here, we take a dynamic programming perspective, [^7] In our numerical results of Section 9, we will take a different, more practical perspective, where we work with a generative model or simulator of the environment. where the model is known and expectations can be computed. It is well known that each iteration of standard VI, which provides a contraction factor of $\gamma$, has time complexity $\mathcal{O}\bigl{(}|\mathcal{S}|^{2}|\mathcal{A}|\bigr{)}$ (Littman et al., 1995). The upper level of FSVI, on the other hand, enjoys an improved contraction factor $\gamma^{T}$ with the same per-iteration running time of $\mathcal{O}\bigl{(}|\mathcal{S}|^{2}|\mathcal{A}|\bigr{)}$, given that we pay a one-time fixed cost of solving the lower level. The quantity $\mathcal{O}\bigl{(}|\mathcal{S}|^{2}|\mathcal{A}|\bigr{)}$ consists of a factor of $|\mathcal{S}||\mathcal{A}|$ due to the number of state-action pairs at which to compute the Bellman update and another factor of $|\mathcal{S}|$ due to the number of successor states. Since freezing slow states restricts the successor states to $\mathcal{Y}$, each iteration of the lower-level VI (first part of Algorithm 2) has running time $\mathcal{O}\bigl{(}|\mathcal{X}||\mathcal{Y}|^{2}|\mathcal{A}|\bigr{)}$. An additional $\mathcal{O}\bigl{(}|\mathcal{S}|^{2}\,T\bigr{)}$ is required to compute the $T$-step transition probabilities of following $\tilde{\bm{\pi}}^{*}$, to be used in the upper-level VI steps, resulting in a one-time fixed cost of $\mathcal{O}(|\mathcal{X}||\mathcal{Y}|^{2}|\mathcal{A}|\,T+|\mathcal{S}|^{2}\,T)$. Particularly when $|\mathcal{X}|$ is large, this can be a reasonable fixed cost to pay in order to get the much improved discount factor of $\gamma^{T}$ going forward. The equations below give a direct comparison between the computational cost of VI vs FSVI:
$\text{Cost}_{\text{VI}}(k)=\;\mathcal{O}\!\bigl{(}k\,|\mathcal{S}|^{2}\,|\mathcal{A}|\bigr{)},$
$\text{Cost}_{\text{FSVI}}(k)=\;\mathcal{O}\!\Bigl{(}\,\underbrace{|\mathcal{X}|\,|\mathcal{Y}|^{2}\,|\mathcal{A}|\,T\;+\;|\mathcal{S}|^{2}\,T}_{\begin{subarray}{c}\text{one-time lower-level}\\ \text{pre\textendash processing cost}\end{subarray}}\;+\;\underbrace{k\,|\mathcal{S}|^{2}\,|\mathcal{A}|}_{\begin{subarray}{c}\text{upper-level}\\ \text{iterations}\end{subarray}}\,\Bigr{)}.$
@@
In this section, we develop the theory to understand the regret of FSVI. First, in Section 6.1, we derive the reward approximation error between the hierarchical reformulation and the frozen-state approximation. This result becomes needed in later sections. In Section 6.2, we define the notion of regret that we use in the paper. Then in Section 6.3, we give an important lemma that illustrates the trade-offs between the various sources of error and how they contribute to the overall regret. Finally in Section 6.4, we give the main result of the section, which characterizes the regret of FSVI for a particular $T$ and $k$.
@@
The detailed proof is in Appendix E.2. ∎
@@
Defining Regret
We start with definitions of the regret of both stationary and $T$-periodic policies.
@@
A General Lemma
In this section, we first prove a general lemma that will be used to analyze FSVI.